OpenStudy (anonymous):

y=x+4 (-7,1) please help me solve this

OpenStudy (anonymous):

just plug in the order pairs

OpenStudy (anonymous):

y=x+4(7) y=x-28 y-28=x OR y=x+4 y+4=x

OpenStudy (anonymous):

(-7, 1) are x,y coordinates. Just plug in the numbers. y=x+4 1=x+4 1-4=x -3= x and to solve for y y=(-7) + 4 y= -3 So the answer is (-3, -3) Hope this helps :)